How many atoms are in 42.4 g Mg ?
1
Expert's answer
2020-05-21T13:52:29-0400

N=n*N(Avogadro)

n(Mg)=m(Mg)/Mr(Mg)=42,4/24=1,77 moll

=>N=1,77*(6,02*10^23)= 1,07*10^24